### v2

# Effective algorithms for tensor-train decomposition via the UTV framework

Abstract
The tensor-train (TT) decomposition is widely used to compress large tensors into a more compact form by exploiting their inherent data structures. A fundamental approach for constructing the TT format is the well-known TT-SVD method, which performs singular value decompositions (SVDs) on the successive matrices sequentially. But in practical applications, it is often unnecessary to compute full SVDs. In this article, we propose a new method called the TT-UTV. It utilizes the virtues of rank-revealing UTV decomposition to compute the TT format for a large-scale tensor, resulting in lower computational cost. We analyze the error bounds on the accuracy of these algorithms in both the URV and ULV cases and then recommend different sweep patterns for these two cases. Based on the theoretical analysis, we also formulate the rank-adaptive algorithms with prescribed accuracy. Numerical experiments on various applications, including magnetic resonance imaging data completion, are performed to illustrate their good performance in practice.
keywords:
Tensor-train decomposition, UTV decomposition , Rank-adaptive algorithms , Data completion.
journal: Nuclear Physics B
[…]
[author1]organization=School of Mathematical Sciences,addressline=Chengdu University of Technology, city=Chengdu, postcode=610059, state=Sichuan, country=China
[…]
[author2]organization=School of Mathematics and Statistics,addressline=Guizhou University, city=Guiyang, postcode=550025, state=Guizhou, country=China
[…]
[author3]organization=School of Mathematical Sciences and Key Laboratory of Mathematics for Nonlinear Sciences,addressline=Fudan University, city=Shanghai, postcode=200433, state=Shanghai, country=China
1 Introduction
Real-world high-dimensional data are usually represented as tensors, such as genetic sequences, video data, hypergraphs, hyperspectral images, and social network interactions [lu2011survey, hore2016tensor, cui2024discrete, velasco2013classification, Koniusz2021, ding2018tensor]. Tensors provide a natural way to model multi-dimensional data and multi-modal nonlinear interactions. However, directly processing these higher-order tensors demands a significant amount of resources, as the time and space complexity grow exponentially with tensor order. This phenomenon is known as the “curse of dimensionality” [Indyk1998]. Fortunately, tensors in the real world often exhibit a low-rank structure, and the tensor-train decomposition [oseledets2011tensor, oseledets2010tt], also known as the matrix product state representation [Perez-Garcia2007MPS], can exploit the inherent low-rank structure to mitigate the curse of dimensionality in many practical scenarios.
The tensor-train (TT) decomposition has found widespread applications in areas such as machine learning [yang2017tensor, novikov2015tensorizing, Sidiropoulos2017], quantum many-body physics [Schollwock2011wave], signal processing [Sedighin2020matrix, miron2020tensor], data completion [xu2023tensor, Steinlechner2016, Cai2022provable], and numerical partial differential equations [Dolgov2021, Richter2021Solving]. For instance, the TT decomposition has been employed to compress deep neural networks while maintaining competitive performance [yang2017tensor, novikov2015tensorizing]. In quantum physics, the TT format enables efficient representation and manipulation of wave functions and operators, thereby significantly reducing computational demands [Schollwock2011wave]. Over the past decade, the TT model has also been adopted for certain high-dimensional data completion problems [Steinlechner2016, Cai2022provable]. In addition, TT-based methods are increasingly applied in solving high-dimensional partial differential equations, where traditional methods often become computationally prohibitive [Dolgov2021, Richter2021Solving], to name a few.
The research on efficient algorithms for constructing TT decomposition of large-scale tensors remains an active area. For a $d$th-order tensor, the TT-SVD algorithm [oseledets2011tensor] sequentially utilizes d1d-1 singular value decompositions (SVDs) to extract core tensors one by one, which serves as a foundational approach to decomposing a higher-order tensor into the TT format. TT-cross approximation [oseledets2010tt] replaces the full SVD computation in the TT-SVD algorithm with skeleton decomposition, making it particularly suitable for large-scale tensors with missing or incomplete data. The alternating least squares method, as an iterative method, computes the TT decomposition by solving a sequence of least squares problems to iteratively refine the TT cores. This method is relatively dependent on the selection of initialization and may exhibit slower convergence for tensors with large condition numbers[Holtz2012the]. In recent years, experts have also proposed parallel algorithms to reduce the runtime for computing tensor-train decomposition; their subroutines still rely on the matrix SVD [zniyed2020tt, shi2023parallel, xie2023error].
In many applications, the SVD is computationally demanding and difficult to update. These limitations have motivated the development of the rank-revealing two-sided orthogonal decomposition, also called the UTV decomposition [Stewart1992an]. The UTV decomposition provides approximately the same numerical rank and subspace information as SVD but at a lower computational cost [Stewart1992an, Fierro1997low, Fierro1995bounding, Demmel2007fast]. Furthermore, UTV decomposition can be efficiently implemented on modern hardware architectures by leveraging block matrix operations that are amenable to parallel processing [Martinsson2019rand]. Due to these virtues, it has become a common alternative to SVD in many applications, including signal processing and machine learning [yang1995projection, Zhou2012stable]. Recently, the UTV decomposition has also been extended to the decompositions of higher-order tensors. For example, a truncated multilinear UTV decomposition [Vandecappelle2022from] is introduced for tensors from the truncated multilinear SVD [DeLathauwer2000a, Vannieuwenhoven2012a]; the t-UTV decomposition [che2022efficient] is proposed for third-order tensors as a substitute for the t-SVD [kilmer2011factorization] based on the tensor t-product.
In this article, we investigate whether the UTV decomposition can be leveraged to construct the TT decomposition as an economic alternative to the TT-SVD. The goal is to save computational costs by utilizing the advantages of UTV decomposition over SVD for large-scale unfolding matrices. This new approach, termed TT-UTV, is designed to operate under both fixed TT-ranks and fixed precision. There exist two types of TT-UTV algorithms corresponding to the URV and ULV cases, respectively. We provide an in-depth error analysis for these algorithms and demonstrate their effectiveness through various numerical experiments, including video recovery and magnetic resonance imaging data completion.
The rest of this paper is organized as follows. Section 2 reviews the necessary preliminary notions. In Section 3, we introduce the TT-ULV algorithm in a left-to-right sweep and TT-URV algorithm in a right-to-left sweep, respectively. We derive rigorous error bounds for them, which explain how the UTV decompositions are used to achieve accurate and efficient TT decomposition, and their computational cost is analyzed. In Section 4, we conduct performance evaluations of the TT-UTV algorithm with TT-SVD in various applications. The conclusions are drawn in Section 5.
2 Preliminaries
Notations. A $d$th-order tensor is a multidimensional array with $d$ free indices. Matrices are second-order tensors, and vectors are first-order tensors. We mainly inherit the usage of notations from the comprehensive works of Kolda and Bader [kolda2009tensor]. The calligraphic capital letters such as $\mathcal{A}$, boldface capital letters such as $\mathbf{A}$, and boldface lowercase letters such as $\mathbf{a}$ are used to denote the tensors, matrices, and vectors, respectively. $\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ denotes the set of all $d$th-order tensors of dimensions $(I_{1},\dots,I_{d})$ over the real number field $\mathbb{R}$, where $I_{k}$ is the dimension of the $k$th mode of the tensor. The entries of $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ are accessed by $\mathcal{A}_{i_{1}i_{2}\cdots i_{d}}$ or $\mathcal{A}(i_{1},i_{2},\dots,i_{d})$ for $1\leq i_{k}\leq I_{k}$. For a tensor $\mathcal{A}$, we denote by $\mathbf{A}_{k}$ its $k$th unfolding matrix.
2.1 Tensor basics
The tensor-train (TT) decomposition represents a higher-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ as the contractions of small factor tensors
$\mathcal{A}_{i_{1}\cdots i_{d}}=\sum_{\alpha_{0}=1}^{r_{0}}\sum_{\alpha_{1}=1}^{r_{1}}\cdots\sum_{\alpha_{d}=1}^{r_{d}}\mathcal{G}^{(1)}_{\alpha_{0}i_{1}\alpha_{1}}\mathcal{G}^{(2)}_{\alpha_{1}i_{2}\alpha_{2}}\cdots\mathcal{G}^{(d)}_{\alpha_{d-1}i_{d}\alpha_{d}}$, (1)
where $r_{0}=r_{d}=1$, these third-order tensors $\mathcal{G}^{(k)}\in\mathbb{C}^{r_{k-1}\times I_{k}\times r_{k}}$ for $k=1,2,\ldots,d$ are called TT-cores, and $(r_{1},r_{2},\ldots,r_{d-1})$ represent TT-ranks. Fig. 1 is a graphical representation of the TT decomposition of a $d$th-order tensor, with core tensors depicted as blue blocks and indices represented as connecting lines. It looks like a train, as its name implies. In virtue of the TT decomposition, the tensor $\mathcal{A}\in\mathbb{C}^{I_{1}\times I_{2}\times\cdots\times I_{k}}$ can be compressed into ${\textstyle\sum_{k=1}^{d}}r_{k-1}I_{k}r_{k}$ parameters, which becomes linearly dependent on the tensor order $d$. Lower TT-ranks imply lower memory consumption and computational costs.
Fig. 1: Graphical illustration of the TT decomposition for a $d$th-order tensor.
Tensor unfolding, the process of rearranging the elements of a tensor into a matrix or vector, is a frequently used operation in tensor analysis and numerical computations [kolda2009tensor]. The reverse lexicographic ordering (as the ‘reshape’ command in MATLAB) is one of the most commonly used approaches for integrating multiple indices into a single long index. Concretely, for the tensor dimensions $\mathbf{I}=\left(I_{1},\dots,I_{d}\right)$ and a corresponding index vector $\mathbf{i}=\left(i_{1},\dots,i_{d}\right)$, the reverse lexicographic order map
$\mathrm{ivec}(\mathbf{i},\mathbf{I}):=i_{1}+\sum_{k=2}^{d}(i_{k}-1)\prod_{j=1}^{k-1}I_{j}.$ (2)
Definition 2.1 (The $k$th unfolding matrix [oseledets2011tensor]).
For a $d$th-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, the matrix $\mathbf{A}_{k}\in\mathbb{R}^{(I_{1}\cdots I_{k})\times(I_{k+1}\cdots I_{d})}$ is called the $k$th unfolding matrix of $\mathcal{A}$ such that
$\mathbf{A}_{k}(i,j)=\mathcal{A}_{i_{1}\cdots i_{k}i_{k+1}\cdots i_{d}},$
where $i=\mathrm{ivec}\left((i_{1},\dots,i_{k}),(I_{1},\dots,I_{k})\right)$ and $j=\mathrm{ivec}\left((i_{k+1},\dots,i_{d}),(I_{k+1},\dots,I_{d})\right)$. That is to say, the first $k$ indices and the last dkd-k indices of $\mathcal{A}$ are fused into the row and column indices of the unfolding matrix $\mathbf{A}_{k}$, respectively, following the reverse lexicographic order $\rm ivec$ (2).
Definition 2.2 (Orthogonal TT-cores [oseledets2011tensor]).
The TT-cores $\{\mathcal{G}^{(k)},k=1,\dots,d\}$ in Eq. (1) are called left-orthogonal if for $k=1,\dots,d-1$,
$\mathbf{G}^{(k)\top}_{2}\mathbf{G}^{(k)}_{2}=\mathbf{I}_{r_{k}},$
where $\mathbf{G}^{(k)}_{2}\in\mathbb{R}^{r_{k-1}I_{k}\times r_{k}}$ are the second unfolding matrices of the TT-cores $\mathcal{G}^{(k)}$. They are right-orthogonal if for $k=2,\dots,d$,
$\mathbf{G}^{(k)}_{1}\mathbf{G}^{(k)\top}_{1}=\mathbf{I}_{r_{k-1}},$
where $\mathbf{G}^{(k)}_{1}\in\mathbb{R}^{r_{k-1}\times I_{k}r_{k}}$ are the first unfolding matrices of the TT-cores $\mathcal{G}^{(k)}$, $\mathbf{I}_{r_{k}}\in\mathbb{R}^{r_{k}\times r_{k}}$ denotes the identity matrix.
The TT-SVD algorithm computes the TT format (1) with left-orthogonal cores by a single left-to-right sweep or the right-orthogonal cores by a right-to-left sweep; more details can be found in [oseledets2011tensor].
Definition 2.3 (Mode product [kolda2009tensor]).
The mode-$k$ product of $\mathcal{X}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ with a matrix $\mathbf{A}\in\mathbb{R}^{J\times I_{k}}$ is an $I_{1}\times\cdots\times I_{k-1}\times J\times I_{k+1}\times\cdots\times I_{d}$ tensor, given by
$\left(\mathcal{X}\times_{k}\mathbf{A}\right)_{i_{1}\cdots i_{k-1}ji_{k}\cdots i_{d}}=\sum_{i_{k}=1}^{I_{k}}\mathcal{X}_{i_{1}\cdots i_{k-1}i_{k}i_{k+1}\cdots i_{d}}\mathbf{A}_{ji_{k}}.$
Apart from the tensor-train decomposition, the Tucker format [tucker1966some] is another compact form for large-scale tensors, which represents a $d$th-order tensor via the mode products of a smaller-scaled core tensor and $d$ factor matrices:
$\mathcal{A}=\mathcal{S}\times_{1}\mathbf{U}_{1}\times_{2}\cdots\times_{d}\mathbf{U}_{d},$ (3)
where $\mathcal{S}\in\mathbb{R}^{r_{1}\times\cdots\times r_{d}}$ and $\mathbf{U}_{k}\in\mathbb{R}^{I_{k}\times r_{k}}$ for $1\leq k\leq d$. Directly from these definitions, we can derive the following proposition.
Proposition 2.1.
Let the tensor $\mathcal{A}$ be given in the Tucker format (3). Then the $k$th unfolding matrices $\mathbf{A}_{k}$ and $\mathbf{S}_{k}$ are related by the matrix equation
$\mathbf{A}_{k}=\left(\mathbf{U}_{k}\otimes\cdots\otimes\mathbf{U}_{1}\right)\mathbf{S}_{k}\left(\mathbf{U}_{d}\otimes\cdots\otimes\mathbf{U}_{k+1}\right)^{\top},$
where $\otimes$ denotes the matrix Kronecker product.
For $\mathbf{A}\in\mathbb{R}^{m\times n}$ and $\mathbf{B}\in\mathbb{R}^{p\times q}$, the Kronecker product $\mathbf{A}\otimes\mathbf{B}$ is the $mp\times nq$ block matrix
$\mathbf{A}\otimes\mathbf{B}=\left[\begin{array}[]{ccc}a_{11}\mathbf{B}&\cdots&a_{1n}\mathbf{B}\\ \vdots&\ddots&\vdots\\ a_{m1}\mathbf{B}&\cdots&a_{mn}\mathbf{B}\end{array}\right].$
Proposition 2.2 ([golub2013matrix]).
Let $\mathbf{A}\in\mathbb{R}^{m\times n},\mathbf{B}\in\mathbb{R}^{p\times q}$, $\mathbf{C}\in\mathbb{R}^{n\times l}$, and $\mathbf{D}\in\mathbb{R}^{q\times r}$. Then $(\mathbf{A}\otimes\mathbf{B})^{\top}=\mathbf{A}^{\top}\otimes\mathbf{B}^{\top}$, and $(\mathbf{A}\otimes\mathbf{B})(\mathbf{C}\otimes\mathbf{D})=(\mathbf{A}\mathbf{C})\otimes(\mathbf{B}\mathbf{D})$.
2.2 UTV decompositions
For the matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$, its UTV decomposition takes the form $\mathbf{A}=\mathbf{UTV}^{\rm\top}$, where $\mathbf{U}\in\mathbb{R}^{m\times n}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ have orthonormal columns, and $\mathbf{T}$ is a triangular matrix. There are two different cases of UTV decomposition corresponding to the structure of the matrix $\mathbf{T}$.
If $\mathbf{T}$ is lower triangular, the decomposition is called ULV decomposition:
$\mathbf{A}=\mathbf{U}\begin{bmatrix}\mathbf{L}_{11}&\mathbf{O}_{r\times(n-r)}\\ \mathbf{L}_{21}&\mathbf{L}_{22}\end{bmatrix}\mathbf{V}^{\top}.$ (4)
Here $\mathbf{L}_{11}$ and $\mathbf{L}_{22}$ are lower triangular, $\mathbf{L}_{21}\in\mathbb{R}^{(m-r)\times r}$ with small entries. Moreover, suppose the numerical rank of $\mathbf{A}$ is $r$, then the ULV decomposition is said to be rank-revealing if
$\sigma_{\rm min}(\mathbf{L}_{11})=O(\sigma_{r})\quad\mathrm{and}\quad\left\|\begin{bmatrix}\mathbf{L}_{21}&\mathbf{L}_{22}\end{bmatrix}\right\|_{2}=O(\sigma_{r+1}).$ (5)
If the middle matrix $\mathbf{T}$ is upper-triangular, then the decomposition called URV decomposition, which takes the form
$\mathbf{A}=\mathbf{U}\begin{bmatrix}\mathbf{R}_{11}&\mathbf{R}_{12}\\ \mathbf{O}_{(m-r)\times r}&\mathbf{R}_{22}\end{bmatrix}\mathbf{V}^{\top},$ (6)
where $\mathbf{R}_{11}\in\mathbb{R}^{r\times r}$, $\mathbf{R}_{22}\in\mathbb{R}^{(m-r)\times(n-r)}$ are upper triangular, $\mathbf{O}$ denotes the zero matrix whose elements are all 0, and $\mathbf{R}_{12}\in\mathbb{R}^{r\times(n-r)}$ contains small entries. The URV decomposition is said to be rank-revealing if
$\sigma_{\rm min}(\mathbf{R}_{11})=O(\sigma_{r})\quad\mathrm{and}\quad\left\|\begin{bmatrix}\mathbf{R}_{12}^{\top}&\mathbf{R}_{22}^{\top}\end{bmatrix}\right\|_{2}=O(\sigma_{r+1}).$ (7)
The smaller the norm of the off-diagonal block, the better the truncated low-rank approximations in Eqs. (4) and (6). In fact, the UTV decomposition of a matrix is not unique, and various matrix decompositions proposed in the literature can be regarded as UTV-type, such as the rank-revealing QR decomposition [CHAN198767], the pivoted QLP decomposition [stewart1999qlp], and the SVD, which is also a special case of UTV decomposition when the middle matrix is diagonal. There are many efficient algorithms for UTV decomposition. Stewart [Stewart1992an] introduced the algorithms for UTV decomposition that are well-suited for high-rank matrices. Fierro etal.et\ al. [fierro1999utv, Fierro1997low] proposed the low-rank UTV algorithms and built a Matlab package named ‘UTV Tools’ for efficient rank-revealing decompositions. In recent years, more fast randomized algorithms have been proposed for UTV decomposition, which are flexible for parallel processing in modern architectures [Martinsson2019rand, Lee2018rankrev, kaloorazi2020efficient]. In many practical applications, they can achieve comparable accuracy at a lower computational cost than classical SVD.
3 The proposed methods
In this section, we propose the TT-UTV algorithms for efficiently computing the approximate TT format by sequentially decomposing the cores via UTV factorizations. We discuss two variants of the algorithm, corresponding to the ULV and URV decompositions, respectively. The feasibility of the proposed methods is rigorously supported by theoretical analysis, based on which we recommend distinct update strategies for each case and further introduce fixed-precision algorithms that can adaptively determine the TT-ranks.
3.1 TT-ULV case
If the required TT-ranks are fixed, the TT-ULV algorithm is shown in Algorithm 1, which computes the approximate TT-format tensor with left-orthogonal cores in a left-to-right sweep.
Algorithm 1 TT-ULV algorithm for fixed TT-ranks (left-to-right sweep)
0: A $d$th-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, and fixed TT-ranks $\{r_{0},r_{1},\dots,r_{d}\}$ with $r_{0}=r_{d}=1$.
0: Left-orthogonal $T$-cores $\{\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}\}$ of the approximation $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$’s.
1: Temporary matrix: $\mathbf{C}=\mathrm{reshape}(\mathcal{A},\left[I_{1},I_{2}\cdots I_{d}\right])$, i.e., the first unfolding matrix $\mathbf{A}_{1}$.
2: for $k=1$ to d1d-1 do
3:  $\mathbf{C}:=\operatorname{reshape}\left(\mathbf{C},\left[r_{k-1}I_{k},I_{k+1}\cdots I_{d}\right]\right)$.
4:  Compute the truncated rank-$r_{k}$ approximation from ULV decomposition $\mathbf{C}=\mathbf{U}_{1}\mathbf{L}_{11}\mathbf{V}_{1}^{\top}+\mathbf{E}$, where $\mathbf{L}_{11}\in\mathbb{R}^{r_{k}\times r_{k}}$ is the leading term of $\mathbf{L}$, and $\mathbf{E}$ denotes the residual part in Eq. (4).
5:  New core: $\mathcal{G}^{(k)}=\operatorname{reshape}\left(\mathbf{U}_{1},\left[r_{k-1},I_{k},r_{k}\right]\right)$.
6:  $\mathbf{C}=\mathbf{L}_{11}\mathbf{V}_{1}^{\top}$.
7: end for
[…]
To facilitate the presentation of error analysis, we introduce some notations. The TT-ULV algorithm 1 adopts the left-to-right sweep (steps 2-7). In the $k$th iteration, step 3 computes a temporary matrix $\mathbf{C}\in\mathbb{R}^{r_{k-1}I_{k}\times I_{k+1}\cdots I_{d}}$; we denote its corresponding (dk+2)(d-k+2)th-order tensor by $\mathcal{A}^{(k)}=\mathrm{reshape}(\mathbf{C},[r_{k-1},I_{k},I_{k+1},\cdots,I_{d}])\in\mathbb{R}^{r_{k-1}\times I_{k}\times\cdots\times I_{d}}$, and the (dk+1)(d-k+1)th-order tensor by $\widetilde{\mathcal{A}}^{(k)}=\mathrm{reshape}($ $\mathbf{C},[r_{k-1}I_{k},I_{k+1},\cdots,I_{d}])\in\mathbb{R}^{r_{k-1}I_{k}\times I_{k+1}\times\cdots\times I_{d}}$. Step 4 computes the column-orthogonal matrix $\mathbf{U}_{1}\in\mathbb{R}^{r_{k-1}I_{k}\times r_{k}}$, denoted by $\mathbf{U}^{(k)}$ in the $k$th iteration. With the notations introduced, the whole process generates temporary tensors $\{\mathcal{A}^{(k)}\}_{k=1}^{d-1}$, $\{\widetilde{\mathcal{A}}^{(k)}\}_{k=1}^{d-1}$, and d1d-1 column-orthogonal matrices $\{\mathbf{U}^{(k)}\}_{k=1}^{d-1}$ that give rise to the TT-cores.
Proposition 3.1.
The unfolding matrices of tensors $\{\mathcal{A}^{(k)}\}_{k=1}^{d-1}$ satisfy the following equations
$\mathbf{A}^{(k+1)}_{l}=\left(\mathbf{I}_{I_{k+l-1}}\otimes\cdots\otimes\mathbf{I}_{I_{k+1}}\otimes\mathbf{U}^{(k)\top}\right)\mathbf{A}^{(k)}_{l+1}$, (8)
for $l=1,\dots,d-k+2$ for all $k=1,\cdots,d-1$, where $\mathbf{A}_{l}^{(k)}$ is the $l$th unfolding matrix of tensor $\mathcal{A}^{(k)}\in\mathbb{R}^{r_{k-1}\times I_{k}\times I_{k+1}\times\cdots\times I_{d}}$.
Proof.
In the $k$th iteration of the TT-ULV algorithm 1, the rank-$r_{k}$ truncated ULV decomposition in step 4 can be written as
$\mathbf{A}_{2}^{(k)}=\mathbf{U}^{(k)}\mathbf{A}_{1}^{(k+1)}+\mathbf{E},$
where $\mathbf{A}_{2}^{(k)}\in\mathbb{R}^{r_{k-1}I_{k}\times I_{k+1}\cdots I_{d}},\mathbf{U}^{(k)}\in\mathbb{R}^{r_{k-1}I_{k}\times r_{k}},\mathbf{A}_{1}^{(k+1)}\in\mathbb{R}^{r_{k}\times I_{k+1}\cdots I_{d}}$. According to the orthogonality $\mathbf{U}^{(k)\top}\mathbf{E}=\mathbf{O}$, we have
$\mathbf{A}_{1}^{(k+1)}=\mathbf{U}^{(k)\top}\mathbf{A}_{2}^{(k)}.$ (9)
From the definitions of the notations $\mathcal{A}^{(k)}$ and $\widetilde{\mathcal{A}}^{(k)}$, the $l$th unfolding matrices of $\widetilde{\mathcal{A}}^{(k)}$ are exactly the $(l+1)$th unfolding matrix of $\mathcal{A}^{(k)}$, i.e., $\mathbf{A}^{(k)}_{l+1}$, for $l=1,\cdots,d-k+1$. Clearly, the above Eq. (9) can be rewritten in tensor format
$\mathcal{A}^{(k+1)}=\widetilde{\mathcal{A}}^{(k)}\times_{1}\mathbf{U}^{(k)\top}\times\mathbf{I}_{I_{k+1}}\times\cdots\times\mathbf{I}_{I_{d}}.$
From the above tensor equation, the Eq. (8) holds by Proposition 2.1. ∎
Lemma 3.1.
Let $m\geq s\geq r$ be positive integers, $\mathbf{Q}\in\mathbb{R}^{s\times m}$ be a row-orthogonal matrix with unit norm rows, $\mathbf{A}\in\mathbb{R}^{m\times n}$, and $\mathbf{B}=\mathbf{Q}\mathbf{A}\in\mathbb{R}^{s\times n}$. Then the rank-r truncated error of the ULV (or URV) decomposition of $\mathbf{B}$ can be smaller than that of $\mathbf{A}$.
Proof.
Suppose a rank-$r$ truncated ULV decomposition of $\mathbf{A}$ is
$\mathbf{A}=\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}+\mathbf{E},$
and the truncated error $\|\mathbf{E}\|_{F}=\varepsilon$, where $\widehat{\mathbf{U}}\in\mathbb{R}^{m\times r}$. Then
$\mathbf{B}=\mathbf{Q}\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}+\mathbf{Q}\mathbf{E}.$
Since $\mathbf{Q}$ has orthogonal rows with unit norm, $\|\mathbf{Q}\|_{2}=1$, then $\|\mathbf{Q}\mathbf{E}\|_{F}\leq\|\mathbf{E}\|_{F}=\varepsilon$. For $\mathbf{Q}\widehat{\mathbf{U}}\in\mathbb{R}^{s\times r}$, we can perform the Gram-Schmidt orthogonalization process on its columns to obtain a decomposition (similar to the QR decomposition) $\mathbf{Q}\widehat{\mathbf{U}}=\widetilde{\mathbf{U}}\mathbf{L}$, where $\widetilde{\mathbf{U}}\in\mathbb{R}^{s\times r}$ has orthogonal columns with unit norm and $\mathbf{L}\in\mathbb{R}^{r\times r}$ is lower-triangular. Then $\widetilde{\mathbf{L}}=\mathbf{L}\widehat{\mathbf{L}}$ is still a lower-triangular matrix. Therefore,
$\mathbf{B}=\widetilde{\mathbf{U}}\widetilde{\mathbf{L}}\widehat{\mathbf{V}}^{\top}+\mathbf{Q}\mathbf{E},$
the leading term $\widetilde{\mathbf{U}}\widetilde{\mathbf{L}}\widehat{\mathbf{V}}^{\top}$ is actually a rank-$r$ truncated ULV decomposition of $\mathbf{B}$, and the residual part $\|QE\|_{F}\leq\varepsilon$. Thus, it finds a rank-$r$ truncated ULV decomposition for $\mathbf{B}$ such that the approximate error is smaller than that of $\mathbf{A}$. For the URV case, the above analysis similarly holds. ∎
Remark 3.1.
According to Proposition 2.2, the Kronecker matrices $\left(\mathbf{I}_{I_{k+l-1}}\otimes\cdots\otimes\mathbf{I}_{I_{k+1}}\otimes\mathbf{U}^{(k)\top}\right)$ in the Eq. (8) are row-orthogonal with unit-norm rows.
Next, the error analysis for the TT-ULV algorithm 1 can be derived as shown in Theorem 3.1 and Corollary 3.1, which can be seen as a generalization of the results for TT-SVD [oseledets2011tensor].
Theorem 3.1 (TT-ULV case).
Given a tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$. Assume that the rank-$r_{k}$ approximation error of its $k$th unfolding matrix $\mathbf{A}_{k}\in\mathbb{R}^{(I_{1}\cdots I_{k})\times(I_{k+1}\cdots I_{d})}$ from the ULV decomposition is $\varepsilon_{k}$, that is,
$\mathbf{A}_{k}=\begin{bmatrix}\mathbf{U}^{(k)}_{1}&\mathbf{U}^{(k)}_{2}\end{bmatrix}\begin{bmatrix}\mathbf{L}^{(k)}_{11}&\mathbf{O}\\ \mathbf{L}^{(k)}_{21}&\mathbf{L}^{(k)}_{22}\end{bmatrix}\begin{bmatrix}\mathbf{V}^{(k)\top}_{1}\\ \mathbf{V}^{(k)\top}_{2}\end{bmatrix}=\mathbf{U}^{(k)}_{1}\mathbf{L}^{(k)}_{11}\mathbf{V}^{(k)\top}_{1}+\mathbf{E}^{(k)},$ (10)
where $\mathbf{L}^{(k)}_{11}\in\mathbb{R}^{r_{k}\times r_{k}}$, $\mathbf{E}^{(k)}=\mathbf{U}^{(k)}_{2}(\mathbf{L}^{(k)}_{21}\mathbf{V}^{(k)\top}_{1}+\mathbf{L}^{(k)}_{22}\mathbf{V}^{(k)\top}_{2})$ and $\|\mathbf{E}^{(k)}\|_{F}=\varepsilon_{k}$ for all $k=1,\dots,d-1$. Then the approximate tensor $\widehat{\mathcal{A}}$ with TT-ranks not higher than $\{r_{k}\}_{0}^{d}$, computed via the TT-ULV algorithm 1 in a left-to-right sweep, meets
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}\leq\sqrt{\sum_{k=1}^{d-1}\varepsilon_{k}^{2}}.$
Proof.
In the TT-ULV algorithm 1, it yields the first TT-core from the ULV decomposition of the first unfolding matrix
$\mathbf{A}_{1}=\mathbf{U}_{1}^{(1)}\mathbf{A}_{2}^{(2)}+\mathbf{E}^{(1)},$
where $\mathbf{U}^{(1)}_{1}\in\mathbb{R}^{I_{1}\times r_{1}}$ leads to the first TT-core, and $\mathbf{A}_{2}^{(2)}=\mathbf{L}^{(1)}_{11}\mathbf{V}^{(1)\top}_{1}\in\mathbb{R}^{r_{1}\times(I_{2}\cdots I_{d})}$ is used to produce the subsequent TT-cores. Note that $\mathbf{A}_{2}^{(2)}$ is the second unfolding matrix of the tensor $\mathcal{A}^{(2)}$. Based on the column orthogonality, it follows that $\mathbf{U}_{1}^{(1)\top}\mathbf{E}^{(1)}=\mathbf{O}$. As we can see, the subsequent processes are to approximate the (d1)(d-1)th-order tensor $\widetilde{\mathcal{A}}^{(2)}=\mathrm{reshape}(\mathbf{A}_{2}^{(2)},[r_{1}I_{2},I_{3},\cdots,I_{d}])$ folded from the matrix $\mathbf{A}_{2}^{(2)}$. We denote by $\mathcal{B}\in\mathbb{R}^{r_{1}I_{2}\times I_{3}\times\cdots\times I_{d}}$ the approximation of $\widetilde{\mathcal{A}}^{(2)}$ computed by the algorithm, and $\mathbf{B}=\mathrm{reshape}(\mathcal{B},[r_{1},(I_{2}\cdots I_{d})])$. Then, the approximate error of the output tensor $\widehat{\mathcal{A}}$ of the TT-ULV algorithm is
$\begin{split}\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}^{2}&=\|\mathbf{A}_{1}-\mathbf{U}_{1}^{(1)}\mathbf{B}\|_{F}^{2}\\ &=\|\mathbf{A}_{1}-\mathbf{U}_{1}^{(1)}\mathbf{A}_{2}^{(2)}+\mathbf{U}_{1}^{(1)}\mathbf{A}_{2}^{(2)}-\mathbf{U}_{1}^{(1)}\mathbf{B}\|_{F}^{2}\\ &\leq\|\mathbf{E}^{(1)}\|_{F}^{2}+\|\mathbf{A}_{2}^{(2)}-\mathbf{B}\|_{F}^{2}\\ &=\varepsilon_{1}^{2}+\|\widetilde{\mathcal{A}}^{(2)}-\mathcal{B}\|_{F}^{2},\end{split}$ (11)
where the third inequality is derived from the orthogonality of $\mathbf{U}_{1}^{(1)\top}\mathbf{E}^{(1)}=\mathbf{O}$ and $\|\mathbf{U}_{1}^{(1)}\|_{2}=1$. Therefore, it reduces to the estimation of the approximate error for the (d1)(d-1)th-order tensor $\widetilde{\mathcal{A}}^{(2)}$ by TT-ULV algorithm. As explained in Proposition 3.1, Lemma 3.1, and Remark 3.1, the truncated error of rank-$r_{k}$ approximation from the ULV decomposition of the (k1)(k-1)th unfolding matrix of $\widetilde{\mathcal{A}}^{(2)}$ can be smaller than the approximate error $\varepsilon_{k}$ of the rank-$r_{k}$ truncated ULV decomposition for $\mathbf{A}_{k}$, for all $k=1,\dots,d-1$. Hence, based on the above analysis and by the method of mathematical induction, we get
$\|\widetilde{\mathcal{A}}_{2}-\mathcal{B}\|_{F}^{2}\leq\sum_{k=2}^{d-1}\varepsilon_{k}^{2}.$
Combine it with the Eq. (11), the theorem is proved. ∎
As established by Theorem 3.1, the approximation error of the TT-ULV algorithm is bounded by the low-rank approximation errors of the ULV decompositions applied to its unfolding matrices. This result also reveals that the low-rankness of these unfolding matrices directly implies a low-rank TT structure. Furthermore, the analysis also allows us to obtain the final approximation error of the TT-UTV algorithm 1, showing how it propagates from the local errors incurred at each step of the sequential truncated UTV decompositions.
Corollary 3.1.
For the TT-ULV algorithm 1 in the left-to-right sweep, if the error of the $k$th truncated ULV decomposition in step 4 is $\varepsilon_{k}$, then the total error of the computed TT-approximation tensor $\widehat{\mathcal{A}}$ satisfies
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}\leq\sqrt{\sum_{k=1}^{d-1}\varepsilon_{k}^{2}}.$
Algorithm 2 TT-ULV algorithm for prescribed accuracy (left-to-right sweep)
0: A $d$th-order tensor $\mathcal{A}$, and prescribed relative error tolerance $\varepsilon$, and the error weights $\{w_{k}\}_{1}^{d-1}$ with $\sum_{k=1}^{d-1}w_{k}^{2}=1$.
0: Left-orthogonal $T$-cores $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}$ of the approximation tensor $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$ not higher than the ULV truncated ranks of the unfolding matrices $\mathbf{A}_{k}$ such that the truncated error $\varepsilon_{k}=w_{k}\varepsilon\|\mathcal{A}\|_{F}$. Then from Corollary 3.1, the computed TT-approximation tensor satisfies
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}\leq\varepsilon\|\mathcal{A}\|_{F}.$
1: Compute truncation parameters $\varepsilon_{k}=w_{k}\varepsilon\|\mathcal{A}\|_{F}$, set $r_{0}=1$.
2: Temporary matrix: $\mathbf{C}=\mathrm{reshape}(\mathcal{A},[I_{1},I_{2}\cdots I_{d}])$.
3: for $k=1$ to d1d-1 do
4:  $\mathbf{C}:=\mathrm{reshape}\left(\mathbf{C},\left[r_{k-1}I_{k},I_{k+1}\cdots I_{d}\right]\right)$.
5:  Compute truncated ULV decomposition $\mathbf{C}=\mathbf{U}_{1}\mathbf{L}_{11}\mathbf{V}_{1}^{\top}+\mathbf{E}$, and determine the rank $r_{k}$ such that $\mathbf{L}_{11}\in\mathbb{R}^{r_{k}\times r_{k}}$ and $\|\mathbf{E}\|_{F}\leq\varepsilon_{k}$.
6:  New core: $\mathcal{G}^{(k)}:=\operatorname{reshape}\left(\mathbf{U}_{1},\left[r_{k-1},I_{k},r_{k}\right]\right)$.
7:  $\mathbf{C}:=\mathbf{L}_{11}\mathbf{V}_{1}^{\top}$.
8: end for
[…]
Therefore, we also present the version of TT-ULV algorithm for prescribed accuracy in Algorithm 2 that can adaptively determine the TT-ranks. In this algorithm, if equal error weights are used, then $w_{k}=\frac{1}{\sqrt{d-1}}$. It is worth noting that in Step 5, we need to determine the smallest rank $r_{k}$ by comparing the truncation residual of the ULV decomposition with the given error tolerance $\varepsilon_{k}$. In the ULV desomposition $\mathbf{C}=\mathbf{U}_{1}\mathbf{L}_{11}\mathbf{V}_{1}^{\top}+\mathbf{E}$, the orthgonolity of $\mathbf{U}_{1}^{\top}\mathbf{E}=\mathbf{O}$ implies that $\|\mathbf{E}\|_{F}^{2}=\|\mathbf{C}\|_{F}^{2}-\|\mathbf{L}_{11}\|_{F}^{2}$. Thereby, we only need to compute the Frobenius norms of $\mathbf{C}$ and the row vectors of $\mathbf{L}_{11}$. Then the truncated rank $r_{k}$ can be determined as the smallest integer for which $\|\mathbf{L}_{11}\|_{F}^{2}\geq\|\mathbf{C}\|_{F}^{2}-\varepsilon_{k}^{2}$. If $\mathbf{C}$ is of size $m\times n$, then this process has a computational complexity of $O(mn)$, which is substantially lower than that of the preceding ULV decomposition.
As we can see, an essential requisite for the above error analysis for Algorithm 1 is that in each iteration the column-orthogonal matrix $\mathbf{U}_{1}$ used to generate the TT-core is orthogonal to the residual part $\mathbf{E}$ of the truncated ULV decomposition. Corollary 3.1 theoretically controls the accumulation of the truncated errors by a sharp bound.
However, if the right-orthogonal TT-cores are needed, the algorithm should adopt a right-to-left sweep, as given in Algorithm 4. Please refer to A, where we explain that this algorithm is not recommended for computing the right-orthogonal cores in the right-to-left sweep, especially for the objective of computing a fixed-precision TT decomposition. Interestingly, the URV case is precisely the opposite, which is more suitable to compute the TT format in a right-to-left sweep.
3.2 TT-URV case
Recall the URV decomposition (6)
$\mathbf{A}=\mathbf{U}_{1}\mathbf{R}_{11}\mathbf{V}_{1}^{\top}+(\mathbf{U}_{1}\mathbf{R}_{12}+\mathbf{U}_{2}\mathbf{R}_{22})\mathbf{V}_{2}^{\top},$
the matrix $\mathbf{V}_{1}$ is orthogonal to the residual part $(\mathbf{U}_{1}\mathbf{R}_{12}+\mathbf{U}_{2}\mathbf{R}_{22})\mathbf{V}_{2}^{\top}$. Wherefore, using the URV decomposition to compute the TT format is more suitable for the right-to-left sweep than for the left-to-right sweep.
We formulate the TT-URV algorithm 3 for fixed TT-ranks to replace Algorithm 4 in a right-to-left sweep. The results similar to Proposition 3.1 and Theorem 3.1 also hold for Algorithm 3.
Algorithm 3 TT-URV algorithm for fixed TT-ranks (right-to-left sweep)
0: A $d$th-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, and fixed TT-ranks $\{r_{0},r_{1},\dots,r_{d}\}$ with $r_{0}=r_{d}=1$.
0: Right-orthogonal $T$-cores $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}$ of the approximation $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$’s.
1: Temporary matrix: $\mathbf{C}=\mathrm{reshape}(\mathcal{A},\left[I_{1}\cdots I_{d-1},I_{d}\right])$, i.e., the (d1)(d-1)th unfolding matrix $\mathbf{A}_{d-1}$.
2: for $k=d$ to 22 do
3:  $\mathbf{C}:=\operatorname{reshape}\left(\mathbf{C},\left[I_{1}\cdots I_{k-1},I_{k}r_{k}\right]\right)$.
4:  Compute the truncated rank-$r_{k-1}$ approximation from URV decomposition $\mathbf{C}=\mathbf{U}_{1}\mathbf{R}_{11}\mathbf{V}_{11}^{\top}+\mathbf{E}$, where $\mathbf{U}_{1}\mathbf{R}_{11}\mathbf{V}_{1}^{\top}$ with $\mathbf{R}_{11}\in\mathbb{R}^{r_{k-1}\times r_{k-1}}$ is the leading term, and $\mathbf{E}$ denotes the residual part in the URV decomposition (6).
5:  New core: $\mathcal{G}^{(k)}=\operatorname{reshape}\left(\mathbf{V}_{1}^{\top},\left[r_{k-1},I_{k},r_{k}\right]\right)$.
6:  $\mathbf{C}=\mathbf{U}_{1}\mathbf{R}_{11}$.
7: end for
[…]
Theorem 3.2 (TT-URV case).
Given a tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$. Assume that the rank-$r_{k}$ approximation error of its $k$th unfolding matrix $\mathbf{A}_{k}\in\mathbb{R}^{(I_{1}\cdots I_{k})\times(I_{k+1}\cdots I_{d})}$ from the URV decomposition is $\varepsilon_{k}$, that is,
$\mathbf{A}_{k}=\begin{bmatrix}\mathbf{U}^{(k)}_{1}&\mathbf{U}^{(k)}_{2}\end{bmatrix}\begin{bmatrix}\mathbf{R}^{(k)}_{11}&\mathbf{R}^{(k)}_{12}\\ \mathbf{O}&\mathbf{R}^{(k)}_{22}\end{bmatrix}\begin{bmatrix}\mathbf{V}^{(k)\top}_{1}\\ \mathbf{V}^{(k)\top}_{2}\end{bmatrix}=\mathbf{U}^{(k)}_{1}\mathbf{R}^{(k)}_{11}\mathbf{V}^{(k)\top}_{1}+\mathbf{E}^{(k)},$ (12)
where $\mathbf{R}^{(k)}_{11}\in\mathbb{R}^{r_{k}\times r_{k}}$, $\mathbf{E}^{(k)}=(\mathbf{U}^{(k)}_{1}\mathbf{R}^{(k)}_{12}+\mathbf{U}^{(k)}_{2}\mathbf{R}^{(k)}_{22})\mathbf{V}^{(k)\top}_{2}$, and $\|\mathbf{E}^{(k)}\|=\varepsilon_{k}$ are defined for all $k=1,\dots,d-1$. Then the approximate tensor $\widehat{\mathcal{A}}$ with TT-ranks not higher than $\{r_{k}\}_{0}^{d}$, computed via the TT-URV algorithm 3 in the right-to-left sweep, meets
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}\leq\sqrt{\sum_{k=1}^{d-1}\varepsilon_{k}^{2}}.$
The TT-URV algorithm for prescribed accuracy can be obtained analogously, referring to Algorithms 2 and 3. It is not repeated here.
Based on the above theoretical analysis, we recommend the TT-ULV algorithm in a left-to-right sweep to compute the left-orthogonal TT-cores and the TT-URV algorithm in a right-to-left sweep to compute the right-orthogonal TT-cores.
Remark 3.2.
In practice, both the ULV and URV decompositions can be flexibly adapted for computing the TT decomposition with left-orthogonal or right-orthogonal TT-cores. Taking TT-URV as an example, while our theoretical analysis indicates that it is better suited for a right-to-left update pattern, one can still use it to compute left-orthogonal TT cores for a given tensor. This is achieved by simply reversing the index order of the input tensor and applying the TT-URV algorithm with a right-to-left sweep to update the computation on the reversed tensor.
3.3 Computational cost
In both the TT-SVD and TT-UTV algorithms, the predominant computational cost in each iteration arises from low-rank matrix approximation, implemented through either SVD or the more general UTV decomposition. This distinction also represents the essential difference between these two approaches. Consequently, comparing their computational complexities boils down to the differences in cost between matrix SVD and UTV decomposition.
The computational procedure for the SVD typically involves two key stages: an initial reduction of the matrix to bidiagonal form, which is a non-iterative process, followed by an iterative QR phase. For an $m\times n$ matrix where $m\geq n$, the SVD demands $O(mn^{2})$ floating-point operations [golub2013matrix]. In contrast, the algorithms for UTV decompositions encompass a family of related methods, including the high-rank UTV updating method [Stewart1992an], the low-rank revealing UTV decompositions [Fierro1997low], and the rank-revealing QR decomposition [CHAN198767, gu1996efficient], etc. Although rank-revealing UTV algorithms share the same asymptotic complexity of $O(mn^{2})$ with the SVD, the scaling coefficient for the dominant $mn^{2}$ term is generally lower [fierro1999utv, golub2013matrix]. This difference is intuitive, as a UTV decomposition imposes fewer structural constraints on its factors than an SVD, leading to potentially lower computational overhead. In addition, UTV algorithms often make better use of fast BLAS routines and are more flexible for parallel processing compared to the SVD, resulting in comparatively faster execution in practical applications [martinsson2020randomized, Martinsson2019rand].
The computational savings of UTV over SVD are amplified in the TT decomposition of a $d$th-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times I_{2}\times\cdots\times I_{d}}$. Since the TT decomposition requires d1d-1 sequential matrix factorizations, the savings achieved by UTV decomposition at each iteration accumulate across the entire hierarchy. As a result, the UTV-based methods enjoy a scaled computational advantage relative to the TT-SVD method, building upon their inherent efficiency at the single-matrix level.
4 Experiments
In this section, we demonstrate the good performance of the TT-UTV algorithms for different applications. There are various available algorithms for rank-revealing UTV decomposition, such as the low-rank UTV algorithm [Fierro1997low] packaged in the ‘UTV tools’ [fierro1999utv, fierro2005utv], which will be used in the numerical tests. The experiments were conducted on a computer equipped with an Apple M4 chip (featuring a 10-core CPU) and 24 GB of RAM, using MATLAB R2025a. In our benchmark studies, the TT-SVD algorithm is executed using the economy-size SVD command svd(A, ‘econ’) to ensure computational efficiency.
4.1 Video data
The experimental dataset was derived from a source video  [^1] https://youtu.be/X9cCgErwqQghttps://youtu.be/X9cCgErwqQg comprising 3,757 frames at a resolution of 2160×3840. Owing to the substantial memory requirements of the raw video, each frame was downsampled to 72×128 pixels using the MATLAB ‘imresize’ command. Hence, the subsequent separation of the RGB channels produced 11,271 images of size 72×128, which form a third-order tensor of dimensions 11271×72×128, denoted by $\mathcal{V}$. To benchmark the proposed algorithms, we evaluated their efficiency and accuracy on the tensor $\mathcal{V}$ under both fixed-rank and given-precision conditions. Let $\widehat{\mathcal{V}}$ denote the tensor reconstructed from the TT formats generated by these algorithms. The reconstruction accuracy is measured by the Relative Squared Error (RSE), defined as
$\mathrm{RSE}=\frac{\|\widehat{\mathcal{V}}-\mathcal{V}\|_{F}}{\|\mathcal{V}\|_{F}}.$
Computational efficiency is benchmarked by the runtime of each algorithm. To mitigate the effects of randomness, the reported runtime is the average elapsed time over 10 independent executions. Note that the TT-SVD and TT-ULV algorithms are executed via a left-to-right sweep. In contrast, the TT-URV algorithm is applied to the tensor with its indices reversed and proceeds via a right-to-left sweep. All of them yield left-orthogonal TT-core tensors.
The performance of these algorithms was evaluated under fixed TT-ranks $(1,R_{1},R_{2},1)$, as reported in Table 1. As we can see, the numerical results demonstrate that the proposed TT-UTV algorithms effectively reduce computational time while achieving satisfactory accuracy comparable to that of TT-SVD. Additionally, the UTV-based algorithms exhibit superior computational efficiency when smaller TT-ranks are required.
We also evaluated the fixed-precision variants of the algorithms under different error tolerance settings, with the error weights $w_{k}$ set uniformly to $1/\sqrt{2}$. As summarized in Table 2, the TT-ranks determined by the TT-UTV algorithms closely match those produced by the TT-SVD algorithm. As expected, larger error tolerances lead to smaller TT-ranks. In all cases, the actual computational errors remain strictly below, yet close to, the specified error tolerances $\varepsilon$, which confirms the effectiveness of our theoretical error bounds derived for the TT-UTV algorithm. Although the TT-ranks determined by the TT-UTV methods are occasionally slightly higher than those from TT-SVD, the TT-UTV variants often achieve higher accuracy as a compensatory benefit. Moreover, the UTV-based algorithms exhibit advantages in computational efficiency in low-rank scenarios.
Table 1: The numerical results under various fixed $T$-ranks.
    Time (s)     RSE
$(R_{1},R_{2})$     $(5,5)$ $(10,10)$ $(15,15)$ $(20,20)$     $(5,5)$ $(10,10)$ $(15,15)$ $(20,20)$
$T$-SVD     86.0 89.7 97.8 101.0     0.1179 0.0938 0.0832 0.0776
$T$-ULV     41.1 50.6 58.6 64.4     0.1179 0.0939 0.0835 0.0782
$T$-URV     39.2 51.5 58.3 64.8     0.1179 0.0939 0.0835 0.0781
Table 2: The numerical results of fixed precision algorihtms
      $\varepsilon=0.09$     $\varepsilon=0.1$     $\varepsilon=0.11$   Method     $(R_{1},R_{2})$ Time (s) RSE     $(R_{1},R_{2})$ Time (s) RSE     $(R_{1},R_{2})$ Time (s) RSE TT-SVD     $(38,6)$ 87.6 0.0885     $(27,4)$ 83.8 0.0995     $(19,3)$ 84.2 0.1064 TT-ULV     $(38,6)$ 86.4 0.0885     $(27,5)$ 69.4 0.0941     $(19,3)$ 59.8 0.1064 TT-URV     $(39,6)$ 87.7 0.0886     $(28,5)$ 71.0 0.0945     $(19,3)$ 60.3 0.1065
4.2 Image compression
In this experiment, we test the performance of TT-UTV algorithms on image compression. Three color images are picked from the Berkeley Segmentation dataset [martin2001database], which are of size $321\times 481$ pixels. In order to facilitate the dimension factorizations, they are resized into $324\times 486$ dimensions using bicubic interpolation via the MATLAB ‘imresize’ command [ko2020fast]. Then each processed image is represented by three $324\times 486$ matrices corresponding to its RGB channels. We first reshape the channel matrices into higher-order tensors of size $18\times 18\times 18\times 27$, and then compute their low-rank approximation using the TT-SVD and TT-UTV algorithms.
Fig. 2 demonstrates the compression performance on three test images under a fixed TT-rank of $(1,15,45,25,1)$. With this configuration, a compressed image requires only $100,035$ parameters, fewer than the $472,392$ parameters of the original. Visually, these compressed images are too similar to distinguish their resolutions and retain most of the structural and textural information from the original images. Furthermore, Fig. 3 reports the Relative Squared Errors (RSEs) under varying TT-ranks of the form $(1,15,r,25,1)$, where $r$ ranges from 40 to 50 by increments of 1. The results indicate that the RSEs achieved by the three algorithms are very close across all tested ranks.
(a) Original images
(b) $T$-SVD compressions
(c) $T$-URV compressions
(d) $T$-ULV compressions
Fig. 2: Three color images from the Berkeley Segmentation Dataset.
Fig. 3: The RSEs of $T$ compressions on these images are very close, and decline with the $T$-rank $r$.
4.3 MRI data completion
When a portion of the information in high-dimensional data is missing or distorted, we can recover the data based on TT-format tensor completion, which can be solved using the Riemannian gradient descent (RGrad) method [Steinlechner2016]. Unlike the traditional gradient descent method in Euclidean spaces, the RGrad method uses the gradients on Riemannian manifolds formed by the fixed-rank TT-format tensors. From an initial TT-format tensor in the Riemannian manifold, the first step in each iteration is computing the Riemannian gradient $\mathcal{RG}(\mathcal{X}_{k})$ at the current point $\mathcal{X}_{k}$. Then compute the estimated tensor $\widehat{\mathcal{X}}_{k+1}=\mathcal{X}_{k}-\alpha_{k}\mathcal{RG}(\mathcal{X}_{k})$, where $\alpha_{k}$ is the step size, but it may not locate in the prescribed Riemannian manifold. Hence, the next step is retracting $\widehat{\mathcal{X}}_{k+1}$ back to the manifold by TT-SVD algorithm to generate the next iteration point $\mathcal{X}_{k+1}$ [Cai2022provable, wang2019tensor]. In this experiment, we explore the effectiveness of the RGrad method while the retraction step in each iteration is substituted by the TT-UTV algorithms.
The experimental results are evaluated by Relative Square Error (RSE) and Peak Signal-to-Noise Ratio (PSNR), which are commonly used in data inpainting tasks [Li2017low, ko2020fast]. Let $\mathcal{M},\widehat{\mathcal{M}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ denote the ground-truth tensor and the recovered tensor, respectively. Then, the PSNR metric is defined as
$\mathrm{PSNR}$ $=10\log_{10}\frac{\mathcal{M}_{max}^{2}}{\frac{1}{\prod_{k=1}^{d}I_{k}}\|\widehat{\mathcal{M}}-\mathcal{M}\|_{F}^{2}},$
where $\mathcal{M}_{max}$ is the maximum value in the ground-truth tensor $\mathcal{M}$. Smaller RSE and larger PSNR indicate better recovery performance.
We use the cubical magnetic resonance imaging (MRI) data  [^2] https://brainweb.bic.mni.mcgill.ca/brainweb/selection_normal.htmlhttps://brainweb.bic.mni.mcgill.ca/brainweb/selection\_normal.html, which is a third-order tensor denoted by $\mathcal{M}$ of size $181\times 217\times 181$. Firstly, we determine the truncated TT-ranks of this tensor for the Riemannian manifold. We set the prescribed RSE $\varepsilon=0.1$ for the TT decomposition via the TT-SVD and TT-UTV algorithms; all of them produce the approximate tensors in the TT format with TT-ranks equal to $(1,55,41,1)$. We choose a slice $\mathcal{M}_{::36}$ from the MRI data shown in Fig. 4. The four subfigures in the first line present the original slice and the compressed ones via various TT decomposition algorithms. As we can see, these compressed slices preserve the most information from the original tensors.
(a) Original slice
(b) $T$-SVD compression
(c) $T$-ULV compression
(d) $T$-URV compression
(e) 70% missing pixels
(f) RGrad method with TT-SVD, PSNR = $30.0277$
(g) RGrad method with $T$-ULV, PSNR = $30.0279$
(h) RGrad method with $T$-URV, PSNR = $30.0274$
Fig. 4: The reconstruction of one slice $\mathcal{M}_{::36}$ of the MRI brain data.
Afterwards, to test the effectiveness of the RGrad method with TT-UTV algorithms, 70% elements of the original MRI data $\mathcal{M}\in\mathbb{R}^{181\times 217\times 181}$ are randomly removed, and we would like to recover them. We perform three experiments using the RGrad method, where the retracting steps are computed by TT-SVD, TT-ULV and TT-URV algorithms, respectively. The second row in Fig. 4 shows the original slice with missing pixels and the recovered slices by RGrad methods with TT-SVD, TT-ULV and TT-URV, respectively. These algorithms have good performance and recover most of the missing data of the original tensor. In addition, Fig. 5 (a) reports that the RSEs of the recovered tensor decline with the iterations, which converge to a level of about the prescribed accuracy $0.1$, and (b) shows that the PSNRs increase in the process of restoration. These curves are too close to distinguish, which illustrates that the RGrad method with TT-UTV is as effective as the Rgrad method with TT-SVD.
(a) The RSEs of the recovered tensor $\widehat{\mathcal{M}}$
(b) The PSNRs of the recovered tensor $\widehat{\mathcal{M}}$
Fig. 5: The RSEs decline and PSNRs increase with the iterations of the RGrad methods.
5 Conclusions
For the computation of tensor-train decomposition, we employed the rank-revealing UTV decomposition to compute low-rank matrix approximations within the computational process, leading to the TT-UTV algorithm. The algorithm provides a unified framework that incorporates a wide range of UTV-type decompositions for computing the TT decomposition, fully leveraging the advantages of UTV decompositions. Under the algorithmic framework, rigorous theoretical analysis establishes a tight error bound associated with the truncated error in each UTV approximation. Based on this analysis, distinct updating strategies are recommended for the ULV and URV variants, and the fixed-precision version of these algorithms are given. Numerical experiments across various application scenarios demonstrate the effectiveness of the proposed methods. For future work, integrating the UTV decomposition into parallel algorithmic frameworks for TT decomposition, along with the development of their randomized variants constitute an interesting research direction.
Appendix A
The TT-ULV algorithm in a right-to-left sweep is given in Algorithm 4. Regarding the size of the matrix $\widehat{\mathbf{L}}$ in step 4, we consider two distinct cases. If the low-rank approximation $\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}$ in step 4 is of size $\widehat{\mathbf{L}}\in\mathbb{R}^{I_{1}\cdots I_{k-1}\times r_{k-1}}$ instead of $\widehat{\mathbf{L}}\in\mathbb{R}^{r_{k-1}\times r_{k-1}}$, that is, the two terms $(\mathbf{U}_{1}\mathbf{L}_{11}+\mathbf{U}_{2}\mathbf{L}_{21})\mathbf{V}^{\top}_{1}$ in the ULV decomposition (4) are retained to make the matrix $\widehat{\mathbf{V}}=\mathbf{V}_{1}$ orthogonal to the residual term $\mathbf{E}=\mathbf{U}_{2}\mathbf{L}_{22}\mathbf{V}^{\top}_{2}$, then results similar to Theorem 3.1 and Corollary 3.1 also hold for the TT-ULV algorithm 4 in the right-to-left sweep. But this approach leads to extra matrix multiplication $\mathbf{U}_{2}\mathbf{L}_{21}$ in each iteration.
Algorithm 4 TT-ULV algorithm for fixed TT-ranks (right-to-left sweep)
0: A $d$th-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, and fixed TT-ranks $\{r_{0},r_{1},\dots,r_{d}\}$ with $r_{0}=r_{d}=1$.
0: Right-orthogonal $T$-cores $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}$ of the approximation $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$’s.
1: Temporary matrix: $\mathbf{C}=\mathrm{reshape}(\mathcal{A},\left[I_{1}\cdots I_{d-1},I_{d}\right])$, i.e., the (d1)(d-1)th unfolding matrix $\mathbf{A}_{d-1}$.
2: for $k=d$ to 22 do
3:  $\mathbf{C}:=\operatorname{reshape}\left(\mathbf{C},\left[I_{1}\cdots I_{k-1},I_{k}r_{k}\right]\right)$.
4:  Compute the truncated rank-$r_{k-1}$ approximation from URV decomposition $\mathbf{C}=\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}+\mathbf{E}$, where $\mathbf{E}$ denotes the residual part in the ULV decomposition (6).
5:  New core: $\mathcal{G}^{(k)}=\operatorname{reshape}\left(\mathbf{V}_{1}^{\top},\left[r_{k-1},I_{k},r_{k}\right]\right)$.
6:  $\mathbf{C}=\mathbf{U}_{1}\mathbf{L}_{11}$.
7: end for
[…]
The theory of rank-revealing UTV decomposition tells us that the term $\mathbf{U}_{2}\mathbf{L}_{21}$ in Eq. (4) can be very minor. Hence, as presented in Algorithm 4, the middle matrix $\widehat{\mathbf{L}}$ of the ULV truncation in step 4 should only retain the $\mathbf{L}_{11}\in\mathbb{R}^{r_{k-1}\times r_{k-1}}$ to reduce the computational cost. In contrast, the sharp bound in Corollary 3.1 cannot be obtained theoretically, since the matrix $\widehat{\mathbf{V}}=\mathbf{V}_{1}$ used to yield the core tensor is no longer orthogonal to the residual part $\mathbf{E}=\mathbf{U}_{2}\mathbf{L}_{21}\mathbf{V}_{1}^{\top}+\mathbf{U}_{2}\mathbf{L}_{22}\mathbf{V}_{2}^{\top}$. In this case, a larger bound can be given by a similar analysis of Theorem 3.1 and the triangle inequality of the Frobenius norm.
Proposition A.1.
For the TT-ULV algorithm 4 in the right-to-left sweep, if the error of the $k$th truncated ULV decomposition in step 4 is $\varepsilon_{k}$, then the total error of the computed TT-approximation tensor $\widehat{\mathcal{A}}$ satisfies
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}\leq\sum_{k=2}^{d}\varepsilon_{k}.$
Based on the above results, the TT-ULV algorithm in a right-to-left sweep can also be designed for a prescribed accuracy $\varepsilon$. If the ULV truncation in step 4 is taken as $\mathbf{U}_{1}\mathbf{L}_{11}\mathbf{V}^{\top}_{1}$, then the equal truncated error of each ULV decomposition should be $\varepsilon_{k}=\frac{\varepsilon}{d-1}\|\mathcal{A}\|_{F}>\frac{\varepsilon}{\sqrt{d-1}}\|\mathcal{A}\|_{F}$, which may generate the TT-cores with much higher ranks than TT-ULV algorithm 2 in the left-to-right sweep. On the other hand, it can also set $\varepsilon_{k}=\frac{\varepsilon}{\sqrt{d-1}}\|\mathcal{A}\|_{F}$, but the middle matrix of ULV truncation $\widehat{\mathbf{L}}\in\mathbb{R}^{I_{1}\cdots I_{k-1}\times r_{k-1}}$ should retain a larger size, which leads to more computations and storage. Wherefore, we do not strongly recommend the TT-ULV algorithm for computing right-orthogonal cores using a right-to-left sweep. On the contrary, the TT-URV case is well-suited for computing the right-orthogonal cores in a right-to-left sweep; see Section 3.2 for details.